OpenStudy (anonymous):

please double check my work: partial fraction decomposition 3x/x^2-6x+9 factor the denominator 3x/(x-3)^2 so we have A/x-3+B/(x-3)^2 clear the denominator and left with 3x=A(x-3)+B 3x=Ax-3A+B so A=3 and B=9 partial fraction decomposition would be 3/x-3+9/(x-3)^2 right?

jhonyy9 (jhonyy9):

you will get 3x ------- (x-3)*2 - so and with what and how you have continued sorry but i not can understanding if you wan this can you write again like one multiply (3/x-3)(x/(x-3)= 3x/(x-3)*2

OpenStudy (anonymous):

You got it right. The work is perfect. If you want there is a short cut to partial fractions, but it doesn't always work. You can plug in values for x once you got your 3x=A(x-3)+B and make systems of equations or choose the x value to make things simplify nicely. For this problem pick x=3 because you will have 3*3=A(3-3)+B and then solve for B and then you can plug it in and find A. This isn't always the best way to do these problems, but sometimes it makes life easier. As you did this problem is the best way to go about partial fractions.